r 2 20 Metia 59 445 Round the naanwers to 3 decimal giaces Metis
Answer #1

To find the relative frequencies, divide each frequency by the total number of individuals in the sample

(b)
Total frequency = 851560 + 451795 + 59445 + 26475 + 11415 = 1400690
Relative frequencies are
First Nation = 851560 / 1400690 = 0.608
Metis = 451795 / 1400690 = 0.323
Inuit = 59445 / 1400690 = 0.042
Other = 26475/ 1400690 = 0.019
Multiple identities = 11415 / 1400690 = 0.008

(c)
Percentage values are
First Nation = 60.8
Metis = 32.3
Inuit = 4.2
Other = 1.9
Multiple identities = 0.8